OpenStudy (anonymous):

For the graph shown, select the statement that best represents the given system of equations. 2y – x = 2 x – 2y = 8 A. consistent and independent B. not enough information C. inconsistent D. coincident

OpenStudy (anonymous):

Substitute the first equation into the second one. (2y-2)-2y=8 2y-2y-2=8 -2=8 Since negative 2 is not equation to eight, this system of equations is inconsistent.